# In-Situ Fine-Tuning of Wildlife Models in IoT-Enabled Camera Traps for Efficient Adaptation

### v3

@@ 1. Introduction
This domain shift challenge fundamentally limits the deployment of ML models on IoT devices. While larger, more robust models can help maintain accuracy across domains, their computational demands often exceed the capabilities of resource-constrained devices. Recent advances in domain generalization through techniques like domain alignment (muandet2013domain, ), meta-learning (li2018learning, ), and ensemble methods (arpit2022ensemble, ; yeo2021robustness, )) typically require increased model capacity to learn and represent broader feature sets that distinguish different domains. This presents a fundamental tension between achieving high generalization performance and ensuring efficient inference on IoT platforms.
@@
Limitation of Existing Approaches. For resource-constrained IoT devices, model fine-tuning offers a promising approach to address domain shifts without requiring the computational overhead of larger, more complex models like those used in domain generalization approaches (muandet2013domain, ; zhou2022domain, ). However, enabling efficient and effective fine-tuning in remote deployments presents several key challenges.
@@ 2. Background and Motivation
Mathematically, let $\mathcal{X}$ be the input (feature) space and $\mathcal{Y}$ be the target (label) space. A domain is defined as a joint distribution $P(X,Y)$ on $\mathcal{X}\times\mathcal{Y}$. In the context of wildlife classification, $\mathcal{X}$ represents images containing animals of interest (also called animal images) while $\mathcal{Y}$ represents the set of animal labels or classes. We have training data $\mathcal{S}=\{(x,y)\}$ sampled from the source domain distribution $P^{\mathcal{S}}(X,Y)$ to prepare a wildlife classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$. Once the wildlife classifier is deployed, it runs on images sampled from the target domain distribution $P^{\mathcal{T}}(X,Y)$, which could be different from $P^{\mathcal{S}}(X,Y))$ due to domain shift.
@@
Figure 1. Spatial and temporal domain shifts in camera trap applications. In particular, Location 2’s camera position shifted over time while Location 1 and 3’s backgrounds show seasonal changes. As we show in Table 1, these domain shift can cause an 9% - 60% drop in wildlife recognition accuracy.
@@
Figure 2 illustrates the benefits of fine-tuning in addressing domain shifts across different model complexities within the EfficientNet family (efficientnet, ). Models ranging from the lightweight EfficientNet-B0 to the more complex EfficientNet-B4 exhibit substantial gaps in their generalization performance when evaluated on images collected from new locations and different time periods (shown by the blue curve). While larger models inherently possess better domain generalization capabilities, smaller models, though more computationally efficient, tend to perform worse on unseen data. However, fine-tuning these models with new samples collected from the new locations (Figure 2(a)) and different time periods (Figure 2(b)) significantly enhances their accuracy, regardless of their computational complexities (shown by the orange curve). This demonstrates the effectiveness of fine-tuning as a method for adapting models to new domains, effectively bridging the performance gap across varying model sizes.
@@ 3. Design of WildFiT
Figure 3. Overview of WildFiT. The system runs a lightweight classification model on the IoT device and maintains accuracy in the presence of domain shifts through Drift-Aware Fine-Tuning, which uses Background Drift Detection and Class Distribution Drift Detection to identify domain shifts, validates their impact in the Drift Validation module, and triggers model adaptation using Background-Aware Data Synthesis.
@@ 3.2. Background-Aware Data Synthesis
The online phase. The online phase generates synthetic training data by intelligently compositing animal objects with target domain backgrounds. Rather than using naive object placement, we leverage domain-specific insights about wildlife behavior and camera trap deployments to improve synthesis quality. Our approach addresses three key challenges in wildlife monitoring: spatial context, social behavior, and temporal patterns. (1) Location Preserving: We preserve the spatial relationship between animals and their environment by positioning objects based on their original locations using bounding box data from MegaDetector V5, as animals tend to appear in physically meaningful positions relative to terrain features. (2) Herd-Aware: We model social dynamics by synthesizing images with multiple objects for herd animals like zebras, reflecting natural group behaviors that impact both appearance and positioning. (3) Time Preserving: We maintain temporal consistency by matching background selection to known activity patterns - for example, ensuring nocturnal animals appear with night-time backgrounds. These techniques collectively improve fine-tuning accuracy by 4.0%, as shown in Section 4.6.
@@ 3.3. Drift Detection
Reliable drift detection in images remains an unsolved problem due to the high dimensionality of image data. Traditional methods often rely on computationally intensive dimensionality reduction techniques to identify appropriate representations for hypothesis testing (rabanser2019failing, ), rendering them unsuitable for resource-constrained IoT devices.
In contrast to the conventional drift detection methods discussed in Section 5, our approach leverages domain-specific observation that domain shifts in camera trap data primarily arise from either background (environment) changes or animal class distribution changes. When such deviations are identified, the Drift Detection module signals the need for Drift Validation and, if necessary, Model Fine-Tuning.
@@ 3.4. Drift Validation
While detecting domain shifts is important, not all shifts necessitate model adaptation — fine-tuning is only warranted when shifts significantly impact classification accuracy. Drift Validation acts as an intelligent gatekeeper, evaluating whether detected shifts will degrade model performance on incoming animal images. Since actual target domain data is unavailable at validation time, the system leverages the Synthesizer to generate representative test samples that reflect current environmental conditions and class distributions.
@@ 4.1. Experiment Settings
Figure 6 shows the complete end-to-end wildlife camera system capable of capturing images and executing the WildFiT pipeline on the Jetson Orin Nano Super. In this example, images from the D2 dataset are replayed to demonstrate system functionality; however, the pipeline also supports direct input from the onboard camera. A demonstration video showcasing the fully operational end-to-end system is available at (wildfit2025demo, ).
Figure 6. Jetson Orin Nano Super with a camera, LCD display, and a power bank. The LCD display is optional.
@@ Methods
Classification Accuracy Results. Table 3 shows exhaustive results comparing our synthesis method against baselines. We observe the following. (1) Compared to no fine-tuning, our synthesis scheme has a massive gain of 20.5% in accuracy on average (last row), indicating the importance of fine-tuning under domain shifts. (2) Compared to computationally efficient approaches like CutMix (yun2019cutmix, ) and MixUp (zhang2017mixup, ), our method has a significant advantage of 7.3% on average, with 20-30% improvements in some locations where the spatial domain shift is large. These generic augmentation methods often strive to augment images in a domain-agnostic way, which can improve model robustness but may not capture the specific characteristics of new deployment environments. (3) Compared to diffusion model-based approaches, our method generally performs better with 3.0% average gain (and in some locations more than 10%) with much less computation. The accuracy gain comes from the fact that diffusion methods like ControlCom introduce artifacts, and Object-Stitch alters both the background and object, resulting in some low fidelity images in the training set.
@@
Figure 7(a) illustrates the accuracy-efficiency tradeoffs achieved by different approaches. For the Periodic Fine-Tuning baselines, we vary the intervals between fine-tuning events, setting N=2,4,6,8,10,14,21N={2,4,6,8,10,14,21} days to achieve different trade-off points. In contrast, our approach adjusts the thresholds for acceptable accuracy degradation within the Drift Validation module, utilizing thresholds thrthr from $6\%$ to $0\%$ in $1\%$ increments to explore various trade-off scenarios.
Overall, our approach achieves the Pareto frontier of fine-tuning frequency and model accuracy under background drift. By comparing Periodic + BDD (orange curve) against Periodic (green curve), we observe that BDD outperforms random sampling (up to 2%), particularly as the number of fine-tuning rounds increases (smaller N). This highlights the effectiveness of BDD in detecting background drift. Comparing our approach with Periodic + BDD, our method achieves higher accuracy (up to 1.5%) for the same number of fine-tuning rounds. Additionally, to reach a specific accuracy, our approach requires fewer fine-tuning rounds (over 4 rounds) highlighting the effectiveness of the Drift Validation module. This reduction in fine-tuning frequency is especially useful in unattended settings, where minimizing computational overhead and preserving battery life are paramount.
@@ 4.4. End-to-End Performance
Results on Accuracy. For model accuracy, we compare the following approaches:
(1) Pseudolabel: This baseline assumes that a few target domain animal images are collected for model fine-tuning. As ground truth labels are unavailable, it uses the IoT model (EfficientNet-B0), trained on the source domain, to generate pseudo-labels for fine-tuning.
(2) GTlabel: This approach assumes that the target domain animal images have ground truth labels. While impractical, it gives us an upper-bound for the Pseudolabel method.
(3) WildFiT:Syn: This approach removes drift-aware fine-tuning for domain adaptation. We start by initializing our background repository with the first 80 samples and perform a single round of model fine-tuning. This baseline addresses the spatial domain shift only.
(4) WildFiT:Syn+BDD: This approach removes only the CDD module. Like WildFiT-Syn, the background repository is initialized with 80 images, but new backgrounds are added by BDD. Fine-tuning is triggered exclusively by the Drift Validation module in response to detected background drift, with no consideration of class distribution shifts.
(5) WildFiT:All: This is the complete WildFiT.
@@ 4.5. Case Study: Longitudinal Performance
(2) WildFiT:Syn: This method is the same as WildFiT:Syn presented in Table 4.
@@ 4.6. Ablation Studies
Impact of Parameters to Fine-Tune. Table 8 summarizes the impact of parameters to fine-tune on accuracy and one-iteration latency, averaged across 27 test locations in the D2 dataset. Fine-tuning the Fully Connected (FC) layers, biases, and Batch Normalization layers achieves an effective balance between accuracy and latency for on-device fine-tuning. The training latency for updating only the FC layers and biases is comparable to our approach, but including Batch Normalization layers adds a 1.6% accuracy improvement. While training only the FC layers offers faster performance, it results in a significant 6.7% accuracy drop compared to our approach.
@@ 5. Related Work
Domain Adaptation. Domain adaptation (DA) focuses on adapting a model trained on the source domain to perform well on a different but related target domain. Typically, DA methods assume the availability of labeled or unlabeled target data for model adaptation, making them a poor fit for wildlife monitoring applications where data collection is labor-intensive. Some DA approaches, known as zero-shot DA, attempt to generalize to unseen target domains by leveraging auxiliary information about the target domain even if they don’t have direct access to target domain data during training. For example, ZDDA (peng2018zero, ) and CoCoGAN (wang2019conditional, ) learn from the task-irrelevant dual-domain pairs, while Poda (fahes2023poda, ) uses natural language descriptions of the target domain. These approaches cannot be applied, as they rely on auxiliary information that is unavailable for wildlife animal classification tasks. Recently proposed methods, including ZoDi (zodi, ), Min-Max (daynight1, ), and CIConv (daynight2, ), target zero-shot adaptation primarily for day-night or weather-induced shifts. However, these methods assume consistent scene layouts, which are unsuitable for wildlife scenarios and incur heavy computational overhead, making them impractical for resource-constrained deployments.

### v4

# WildFiT: Autonomous In-situ Model Adaptation for Resource-Constrained IoT Systems

@@ 1. Introduction
This domain shift limits the deployment of ML models on IoT devices. While larger, more robust models can help maintain accuracy across domains, their computational demands often exceed the capabilities of resource-constrained devices. Recent advances in domain generalization through techniques like domain alignment (muandet2013domain, ), meta-learning (li2018learning, ), and ensemble methods (arpit2022ensemble, ; yeo2021robustness, )) typically require increased model capacity to learn and represent broader feature sets that distinguish different domains. This presents a fundamental tension between achieving high generalization performance and ensuring efficient inference on IoT platforms.
@@
Limitation of Existing Approaches. For resource-constrained IoT devices, model fine-tuning offers a promising approach to address domain shifts without requiring the computational overhead of larger, more complex models like those used for domain generalization (muandet2013domain, ; zhou2022domain, ). However, enabling efficient and effective fine-tuning in remote deployments presents several key challenges.
@@ 2. Background and Motivation
Figure 1. Spatial and temporal domain shifts in camera trap applications. Specifically, Location 2’s camera position shifted over time while Location 1 and 3’s backgrounds show seasonal changes. As we show in Table 1, these domain shift can cause an 9% - 60% drop in wildlife recognition accuracy.
@@
Figure 2 illustrates the benefits of fine-tuning in addressing domain shifts across different model complexities within the EfficientNet family (efficientnet, ). Models ranging from the lightweight EfficientNet-B0 to the more complex EfficientNet-B4 exhibit substantial gaps in their generalization performance when evaluated on images collected from new locations and different time periods (shown by the blue curve). While larger models inherently possess better domain generalization capabilities, smaller models, though more computationally efficient, tend to perform worse on unseen data. However, fine-tuning these models with new samples collected from the new locations (Figure 2(a)) and different time periods (Figure 2(b)) significantly enhances their accuracy, regardless of their computational complexities (shown by the orange curve). This demonstrates the effectiveness of fine-tuning as a method for adapting models to new domains, bridging the performance gap across varying model sizes.
@@ 3. Design of WildFiT
Figure 3. Overview of WildFiT. The system runs a lightweight classification model on the IoT device and maintains accuracy under domain shifts through Drift-Aware Fine-Tuning, which uses Background Drift Detection and Class Distribution Drift Detection to identify domain shifts, validates their impact in the Drift Validation module, and triggers on-device model adaptation using Background-Aware Data Synthesis.
@@ 3.2. Background-Aware Data Synthesis
The online phase. The online phase generates synthetic training data by intelligently compositing animal objects with target domain backgrounds. Rather than using naive object placement, we leverage domain-specific insights about wildlife behavior and camera trap deployments to improve synthesis quality. Our approach addresses three key challenges in wildlife monitoring: spatial context, social behavior, and temporal patterns. (1) Location Preserving: We retain MegaDetector V5 bounding boxes from source-domain images and place animals at the same relative positions on target-domain backgrounds, as their locations are typically tied to terrain features and camera setup. (2) Herd-Aware: When source images contain multiple animals (herds), we preserve these group configurations including all bounding boxes. During synthesis, we sample random subsets from these groups to create realistic herd compositions while maintaining natural spatial relationships, effectively modeling social dynamics. (3) Time Preserving: We match animal activities with appropriate temporal contexts by pairing diurnal/nocturnal species with corresponding day/night backgrounds. Both the source and target domain data include timestamps, which we leverage to match species with background images captured at corresponding times. These techniques collectively improve fine-tuning accuracy by 4.0%, as shown in Section 4.6.
@@ 3.3. Drift Detection
Reliable drift detection in images remains an unsolved problem due to the high dimensionality of image data. Traditional methods rely on computationally intensive dimensionality reduction techniques to identify appropriate representations for hypothesis testing (rabanser2019failing, ), rendering them unsuitable for resource-constrained IoT devices.
In contrast to the conventional drift detection methods discussed in Section 5, our approach leverages domain-specific observation that domain shifts in camera trap data arise from either background (environment) changes or animal class distribution changes. When such deviations are identified, the Drift Detection module signals the need for Drift Validation and, if necessary, Model Fine-Tuning.
@@ 3.4. Drift Validation
While detecting domain shifts is important, not all shifts necessitate model adaptation — fine-tuning is only warranted when shifts significantly impact classification accuracy. Drift Validation acts as an intelligent gatekeeper, evaluating whether detected shifts will degrade model performance on incoming animal images. Since target domain data is unavailable at validation time, the system leverages the Synthesizer to generate representative test samples that reflect current environmental conditions and class distributions.
@@ Methods
Classification Accuracy Results. Table 3 compares our synthesis method against baselines. We observe the following. (1) Compared to no fine-tuning, our synthesis scheme has a massive gain of 20.5% in accuracy on average (last row), indicating the importance of fine-tuning under domain shifts. (2) Compared to computationally efficient approaches like CutMix (yun2019cutmix, ) and MixUp (zhang2017mixup, ), our method has a significant advantage of 7.3% on average, with 20-30% improvements in some locations where the spatial domain shift is large. These generic augmentation methods often strive to augment images in a domain-agnostic way, which can improve model robustness but may not capture the specific characteristics of new deployment environments. (3) Compared to diffusion-based methods, our approach achieves a 3% average accuracy gain (over 10% in some cases) with much lower computation. This improvement arises because diffusion models often fail to preserve backgrounds and sometimes produce unusable images. Moreover, their slow generation speed prevents on-the-fly synthesis during training, forcing offline image composition. In contrast, our lightweight method supports on-the-fly synthesis for every batch alongside common data augmentation techniques, producing diverse variations (e.g., different herd images) that enhance fine-tuning.
@@
Figure 6(a) illustrates the accuracy-efficiency tradeoffs achieved by different approaches. For the Periodic Fine-Tuning baselines, we vary the intervals between fine-tuning events, setting N=2,4,6,8,10,14,21N={2,4,6,8,10,14,21} days to achieve different trade-off points. In contrast, our approach adjusts the thresholds for acceptable accuracy degradation within the Drift Validation module, utilizing thresholds thrthr from $6\%$ to $0\%$ in $1\%$ increments to explore trade-off scenarios.
Overall, our approach achieves the Pareto frontier of fine-tuning frequency and model accuracy under background drift. By comparing Periodic + BDD (orange curve) against Periodic (green curve), we observe that BDD outperforms random sampling (up to 2%), particularly as the number of fine-tuning rounds increases (smaller N). This highlights the effectiveness of BDD in detecting background drift. Comparing our approach with Periodic + BDD, our method achieves higher accuracy (up to 1.5%) for the same number of fine-tuning rounds. To reach a specific accuracy, our approach requires fewer fine-tuning rounds (over 4 rounds) highlighting the effectiveness of the Drift Validation module. This reduction in fine-tuning frequency is especially useful in unattended settings, where minimizing computational overhead and preserving battery life are paramount.
@@ 4.4. End-to-End Performance
Results on Accuracy. For model accuracy, we compare the following approaches: (1) Pseudolabel: This baseline assumes that a few target domain animal images are collected for model fine-tuning. As ground truth labels are unavailable, it uses the IoT model (EfficientNet-B0), trained on the source domain, to generate pseudo-labels for fine-tuning, following a Test-Time Adaptation (TTA) approach. (2) GTlabel: This approach assumes that the target domain animal images have ground truth labels. While impractical, it gives us an upper-bound for the Pseudolabel method. (3) WildFiT:Syn: This approach removes drift-aware fine-tuning for domain adaptation. We start by initializing our background repository with the first 80 samples and perform a single round of fine-tuning. This baseline addresses the spatial domain shift only. (4) WildFiT:Syn+BDD: This approach removes only the CDD module. Like WildFiT-Syn, the background repository is initialized with 80 images, but new backgrounds are added by BDD. Fine-tuning is triggered by the Drift Validation module in response to detected background drift, with no consideration of class distribution shifts, and (5) WildFiT:All: This is the complete WildFiT.
@@ 4.5. Case Study: Longitudinal Performance
(2) WildFiT:Syn: This method is the same as WildFiT:Syn (Table 4).
@@ 4.6. Ablation Studies
Impact of Parameters to Fine-Tune. Table 8 summarizes the impact of parameters to fine-tune on accuracy and one-iteration latency, averaged across 27 test locations in the D2 dataset. Fine-tuning the Fully Connected (FC) layers, biases, and Batch Normalization (BN) layers achieves an effective balance between accuracy and latency for on-device fine-tuning. In particular, updating the BN layers is inspired by techniques from the Test-Time Adaptation (TTA) literature. The training latency for updating only the FC layers and biases is comparable to our approach, but including Batch Normalization layers adds a 1.6% accuracy improvement. While training only the FC layers offers faster performance, it results in a significant 6.7% accuracy drop compared to our approach.
@@ 5. Related Work
Domain Adaptation. Domain adaptation (DA) focuses on adapting a model trained on the source domain to perform well on a different but related target domain. Typically, DA methods assume the availability of labeled or unlabeled target data for model adaptation, making them a poor fit for wildlife monitoring applications where data collection is labor-intensive. Some DA approaches, known as zero-shot DA, attempt to generalize to unseen target domains by leveraging auxiliary information about the target domain even if they don’t have direct access to target domain data during training. For example, ZDDA (peng2018zero, ) and CoCoGAN (wang2019conditional, ) learn from the task-irrelevant dual-domain pairs, while Poda (fahes2023poda, ) uses natural language descriptions of the target domain. These approaches cannot be applied, as they rely on auxiliary information that is unavailable for wildlife animal classification tasks. Recently proposed methods, including ZoDi (zodi, ), Min-Max (daynight1, ), and CIConv (daynight2, ), target zero-shot adaptation for day-night or weather-induced shifts. However, these methods assume consistent scene layouts, which are unsuitable for wildlife scenarios and incur heavy computational overhead, making them impractical for resource-constrained deployments.
@@
6. Discussion
Test-Time Adaptation (TTA). TTA (tta1, ; SHOT, ; tta3, ; tent, ) enables a model to adjust to distribution shifts during inference by adapting a pre-trained source model to unlabeled target data before making predictions. The key challenge in TTA is achieving robust adaptation with limited or no labeled target samples, a significant problem in applications like wildlife monitoring where data collection is expensive and labels are scarce. Existing TTA methods address this challenge through several strategies, each with distinct trade-offs. Entropy minimization approaches (tent, ) encourage confident predictions by reducing output entropy on test samples, but become unstable when target data is limited. Normalization-based methods (AdaBN, ) realign feature distributions by updating batch normalization layers, offering computational efficiency but only correcting shallow, feature-level shifts that fail under strong domain changes. Pseudo-labeling approaches (SHOT, ) generate labels for target samples to guide self-training, yet reinforce errors when initial predictions are uncertain. We demonstrate this in our evaluation, where the Pseudolabel baseline (a TTA approach) performs poorly. More recent continual TTA methods (CoTTA, ) adapt sequentially in non-stationary environments, but require maintaining multiple models (e.g., student and teacher) and remain prone to catastrophic forgetting. Our synthesis-driven approach addresses TTA’s core limitation i.e. dependence on scarce target data, by generating high-quality, context-aware training samples that capture current environmental conditions. This enables proactive adaptation: the model updates before performance degradation occurs rather than reacting to observed failures. Among our baselines, Pseudolabel represents a standard TTA approach, while GTlabel provides an upper bound with ground-truth labels. Our results show synthesis-based adaptation substantially outperforms these approaches, particularly in data-scarce scenarios.
Unseen species. Our modular design naturally extends to novel or unseen species through two mechanisms. First, novel species detection can leverage out-of-distribution detection (ensemble1, ; mcdrop1, ; maxsoft1, ; testtime2, ; naff1, ; aff2, ) or open-set recognition (openset1, ; openset2, ), together with human review (openset_human, ) or leveraging vision-language models (VLMs). Second, once new species data becomes available, the synthesis pipeline can integrate new animal objects, while fine-tuning can expand the classifier to new classes using class-incremental or few-shot learning. Prior work (bias, ; FC, ; BN, ) shows that fine-tuning limited parameters (e.g., FC, bias, or BN layers) effectively supports such updates without catastrophic forgetting.